OpenStudy (anonymous):

Assume a general exponential growth/decay model. A bacteria culture starts with 500 bacteria and doubles every 3.5 hours. Find an expression for the number of bacteria after t hours.

OpenStudy (anonymous):

550e^35t 500=Qoe^(ln2)/(3.5)(t) 500e^(ln2)/(3.5)(t) 2e^3.5t

OpenStudy (anonymous):

once simple way to do it is to write \[500\times 2^{\frac{t}{3.5}}\] but i guess that is not an option

OpenStudy (anonymous):

you want to solve for \(r\) in the formula \[500e^{rt}\] and you know if \(t=3.5\) then you double to \(1000\) set \[1000=500e^{3.5r}\] and solve for \(r\) via \[2=e^{3.5r}\] \[\ln(2)=3.5r\] \[r=\frac{\ln(2)}{3.5}\]

OpenStudy (anonymous):

therefore you formula is the same as choice C it is always C